# UltraBones100k: A reliable automated labeling method and large-scale dataset for ultrasound-based bone surface extraction

### v2

@@ 3.2 Bone Segmentation
Segmentation Mode. In our study, 14 cadavers were scanned, resulting a dataset of over 100k ultrasound images with bone labels. We split the dataset into 10 training specimens, 1 validation specimen, and 3 test specimens (around 25k ultrasound images). The training and validation sets are further split using cross-validation. A U-net with ResNet encoders is trained on UltraBones100k to automatically segment bones in ultrasound images[32, 33]. The loss function is a weighted combination of Dice Loss (weight=1) and Binary Cross Entropy Loss (weight=1). We trained the model using the Adam optimizer with $\beta=(0.9,0.999)$, starting with an initial learning rate of 1e-5, which decayed by 0.9 every 10 epochs. The training process ran for 100 epochs on an NVIDIA V100 GPU, with a batch size of 32. The code and pretrained model will be publicly available.
@@
The accuracy, completeness, and F1 score with a distance threshold $\sigma$ are defined as follows:
$Acc(\sigma)=\frac{1}{|\mathcal{P}|}\sum_{\hat{\mathbf{x}}\in\mathcal{P}}[d_{\hat{\mathbf{x}}\rightarrow\mathcal{G}}<\sigma]$ (4)
$Com(\sigma)=\frac{1}{|\mathcal{G}|}\sum_{\mathbf{x}\in\mathcal{G}}[d_{\mathbf{x}\rightarrow\mathcal{P}}<\sigma]$ (5)
$F(\sigma)=\frac{2Acc(\sigma)\cdot Com(\sigma)}{Acc(\sigma)+Com(\sigma)},$ (6)
where [.][.][ . ] is the Iverson bracket.
@@ 4 Results
Figure 7: Results of the rating experiment.
Clinical Evaluation. The results of the bone annotation rating experiment is shown in Figure 7. The mean ratings are 1.45 for the initial alignment, 0.87 for CT-US registration, and 2.39 for our optimized alignment. A Wilcoxon signed-rank test with Bonferroni correction demonstrated a significant improvement in bone labeling quality following our optimization pipeline ($p<0.001$).
@@
Generalizability. Our dataset consists of images collected from the lower extremities of human subjects. To qualitatively assess the generalizability of the trained model to new anatomies and specimens, we apply it to ultrasound images of the pelvis and femur collected from new specimens.Qualitative results of the trained model’s generalizability on new anatomies are shown in Figure 9
@@ 5 Discussion
The results highlight our methodology as a promising solution for collecting large-scale ultrasound datasets with automatically generated bone labels. Clinical evaluation results confirm the quality of these labels. Furthermore, the pelvis and hip experiment results demonstrate the effectiveness of the collected dataset in training deep learning models with strong generalizability.
Comparison with SOTA. Previous work [12] aimed to align a tracked CT bone model with tracked ultrasound images. However, they reported that the mismatch of 1–2 mm was too large to allow for accurate automatic bone labeling in ultrasound images, and manual labeling was adopted in the end. In contrast, the results show that our optimization pipeline enhances CT-ultrasound alignment, enabling automatic accurate bone label generation. In addition, unlike the traditional CT-US registration method [23], which calculates a single global transformation matrix for the entire ultrasound sweep to maximize overlapping voxel intensities, our approach computes a transformation matrix for each individual ultrasound frame. According to the results of the clinical evaluation, our method achieves significantly improved bone annotations. This improvement is statistically validated by the Wilcoxon signed-rank test with Bonferroni correction ($p<0.001$).
Automatic vs Manual Labeling. Qualitative results in Figure 6 indicate that the model trained on UltraBones100k outperforms the surgeon in segmenting bones in ultrasound images. Manual labels mostly include only high-intensity bone regions. In contrast, predicted bone labels include both high-intensity and low-intensity regions. Quantitative results in Figure 8 demonstrate that the trained model consistently outperforms surgeon’s manual labeling, regardless of the distance threshold used. These qualitative and quantitative results in combination with the intensity distribution of labels illustrated in Figure 6(b) further confirm that the expert surgeon faces difficulty in correctly labeling bone contours in low-intensity regions. These findings underscore the superiority of our approach over the manual labeling.
@@
Limitations. Although the experiment results demonstrate the generalizability of the model trained on UltraBones100k, its performance may decline in anatomies with significantly different bone-soft tissue structures and shapes, such as the spine and skull. Depending on the required level of accuracy, additional data collection may be necessary for specific anatomies and pathological conditions. In some cases, this may involve the collection of in vivo data. Conducting data collection studies with a limited number of patients is a common and justifiable approach in such scenarios. Furthermore, the publicly available model pretrained on UltraBones100k can be fine-tuned with newly collected data to enhance its performance for these specific cases.
@@ 6 Conclusion
Our methodology presents a promising solution to create large-scale and highly accurate training data for bone surface extraction in ultrasound images without relying on expert labeling. Serving as a benchmark, the segmentation model trained on UltraBones100k outperformed an expert surgeon with significantly more complete bone labels, demonstrating its potential for downstream tasks, such as 3D bone surface reconstruction and registration, and clinical application.
@@ Acknowledgement
This research has been funded by the Innosuisse Flagship project PROFICIENCY No. PFFS-21-19. This work has also been supported by the OR-X, a Swiss national research infrastructure for translational surgery, and associated funding by the University of Zurich and University Hospital Balgrist.

### v3

@@ 3.2 Bone Segmentation
Segmentation Mode. In our study, 14 cadavers were scanned, resulting a dataset of over 100k ultrasound images with bone labels. We split the dataset into 10 training specimens, 1 validation specimen, and 3 test specimens (around 25k ultrasound images). The training and validation sets are further split using cross-validation. A U-net with ResNet encoders is trained on UltraBones100k to automatically segment bones in ultrasound images[33, 34]. The loss function is a weighted combination of Dice Loss (weight=1) and Binary Cross Entropy Loss (weight=1). We trained the model using the Adam optimizer with $\beta=(0.9,0.999)$, starting with an initial learning rate of 1e-5, which decayed by 0.9 every 10 epochs. The training process ran for 100 epochs on an NVIDIA V100 GPU, with a batch size of 32. The code and pretrained model are available at https://github.com/luohwu/UltraBones100k.git.
@@
The accuracy, completeness, and F1 score with a distance threshold $\sigma$ are defined using Iverson brackets [.][.][ . ], where $[Y]$ equals 1 if condition $Y$ is true, and 0 otherwise:
$Acc(\sigma)=\frac{1}{|\mathcal{P}|}\sum_{\hat{\mathbf{x}}\in\mathcal{P}}[d_{\hat{\mathbf{x}}\rightarrow\mathcal{G}}<\sigma],$ (4)
$Com(\sigma)=\frac{1}{|\mathcal{G}|}\sum_{\mathbf{x}\in\mathcal{G}}[d_{\mathbf{x}\rightarrow\mathcal{P}}<\sigma],$ (5)
$F(\sigma)=\frac{2Acc(\sigma)\cdot Com(\sigma)}{Acc(\sigma)+Com(\sigma)}.$ (6)
@@ 4 Results
Clinical Evaluation. The results of the bone annotation rating experiment are presented in Table 2. For rater 1, the mean ratings are 1.45 for the initial alignment, 0.87 for CT-US registration, and 2.39 for our optimized alignment. For rater 2, the corresponding mean ratings are 1.83, 1.15, and 2.56. Wilcoxon signed-rank tests with Bonferroni correction indicate a significant improvement in bone labeling quality following our optimization pipeline ($p<0.001$) for both raters.
Rater ID Initial Labeling Global CT-US Registration Ours
Rater 1 $1.45\pm 0.66$ $0.87\pm 0.71$ $2.39\pm 0.76$
Rater 2 $1.83\pm 0.67$ $1.15\pm 0.81$ $2.56\pm 0.62$
Table 2: Results of the rating experiment. Each cell reports the mean value along with the standard deviation.
@@
Generalizability. Our dataset consists of images collected from the lower extremities of human subjects. To qualitatively assess the generalizability of the trained model to new anatomies and specimens, we apply it to ultrasound images of the pelvis and femur collected from new ex-vivo specimens. Qualitative results of the trained model’s generalizability on ex-vivo anatomies are shown in Figure 8. The model trained UltraBones100k successfully segment the bone surface in ultrasound images acquired from new bones, covering both the high intensity regions and low intensity regions. In addition, to qualitatively assess the model’s generalizability on in-vivo ultrasound data, we apply it to in-vivo ultrasound images of the fibula and tibia. The results in Figure 9 demonstrate that the model can accurately segment the bone surface in in-vivo ultrasound images, despite being trained exclusively on ex-vivo data.
@@
Figure 9: Examples of ultrasound images depicting the in-vivo fibula and tibia, along with the corresponding predicted bone labels.
@@ 5 Discussion
The results highlight our methodology as a promising solution for collecting large-scale ultrasound datasets with automatically generated bone labels. Clinical evaluation results confirm the quality of these labels. Furthermore, the femur and hip experiment results demonstrate the effectiveness of the collected dataset in training deep learning models with strong generalizability.
Comparison with SOTA. Previous work [12] aimed to align a tracked CT bone model with tracked ultrasound images. However, they reported that the mismatch of 1–2 mm was too large to allow for accurate automatic bone labeling in ultrasound images, and manual labeling was adopted in the end. In contrast, the results show that our optimization pipeline enhances CT-ultrasound alignment, enabling automatic accurate bone label generation. In addition, the applied global CT-US registration computes a single transformation matrix to align the original CT volume with the reconstructed ultrasound volume based on features such as intensity [23]. In the original CT volume, we observed local regions with high feature values, which can cause the optimization process to converge to local minima. As a result, the labeling quality becomes inconsistent: within the same ultrasound volume, some frames have high-quality labeling, while others perform poorly. Instead, our approach computes a transformation matrix for each individual ultrasound frame. According to the results of the clinical evaluation, our method achieves significantly improved bone annotations. This improvement is statistically validated by the Wilcoxon signed-rank test with Bonferroni correction ($p<0.001$). However, our approach is significantly more computationally expensive and relies heavily on accurate local initialization from the initial tracking data.
Automatic vs Manual Labeling. Qualitative results in Figure 6 indicate that the model trained on UltraBones100k outperforms the surgeon in segmenting bones in ultrasound images. Manual labels mostly include only high-intensity bone regions. In contrast, predicted bone labels include both high-intensity and low-intensity regions. Quantitative results in Figure 7 demonstrate that the trained model consistently outperforms surgeon’s manual labeling, regardless of the distance threshold used. Note that, compared to the trained model, manual labeling achieves comparable performance in terms of accuracy within the high-intensity regions. This result aligns well with our hypothesis that surgeons can accurately label bone surfaces in these regions. These qualitative and quantitative results in combination with the intensity distribution of labels illustrated in Figure 6(b) further confirm that the expert surgeon faces difficulty in correctly labeling bone contours in low-intensity regions. These findings underscore the superiority of our approach over the manual labeling.
@@
Limitations. Although the experiment results demonstrate the generalizability of the model trained on UltraBones100k, its performance may decline in anatomies with significantly different bone-soft tissue structures and shapes, such as the spine and skull. In addition, UltraBones100k consists of data only acquired from ex-vivo specimens. Compared to in-vivo specimens, the properties of biological tissues, such as elasticity and acoustic characteristics, change in cadaveric specimens, which can result in difference in ultrasound imaging. The distinction between ex-vivo and in-vivo ultrasound data is not a concern in scenarios that exclusively involve ex-vivo specimens, such as hands-on surgical skill training using cadaveric specimens [36]. However, for in-vivo applications, the difference can lead to a declined model performance. Although the results in Figure 9 show that a model trained exclusively on ex-vivo data can accurately segment bone surfaces in in-vivo ultrasound images, additional in-vivo data collection may be necessary for fine-tuning the model [37, 38], depending on the required level of accuracy. Conducting data collection studies with a limited number of patients is a common and justifiable approach in such scenarios. Furthermore, the publicly available model pretrained on UltraBones100k can be fine-tuned with newly collected data to enhance its performance for these specific cases. In addition, the segmentation performance can potentially be improved with network design and architecture optimization [39, 40, 41], particularly by considering the imaging characteristics of ultrasound. Although manual labeling is costly, obtaining annotations from multiple experts is valuable for comparison with automatic labeling and for evaluating both intra-rater and inter-rater variability.
@@ 6 Conclusion
Our methodology presents a promising solution to create large-scale and highly accurate training data for bone surface extraction in ultrasound images without relying on expert labeling. Serving as a benchmark, the segmentation model trained on UltraBones100k outperformed an expert surgeon with significantly more complete bone labels. The methodology and dataset show promise in advancing research and clinical translation of ultrasound imaging in computer-assisted interventions, such as surgical planning and intraoperative guidance, where 2D bone segmentation, 3D bone surface reconstruction, and multi-modality bone registration play critical roles.
@@ Acknowledgement
This research has been funded by the Innosuisse Flagship project PROFICIENCY No. PFFS-21-19. This work has also been supported by the OR-X, a Swiss national research infrastructure for translational surgery, and associated funding by the University of Zurich and University Hospital Balgrist. We appreciate Andreas Schweizer for his support and help.